Problem is:
Fixed cost, $200; 50 items cost $2000 to produce - what is appropriate cost function?
The answer is c(x) = 36x + 200 Can you please explain how they got to that answer. Thank you.

Answer by tutorcecilia(2152)

y=mx+b [start with the basic equation of a line]
(x, y) = (50, 2000) [identify the given variables]
b= fixed cost = 200 [identify the given variables]
.
y=mx+b
2000=m(50)+200 [plug-in the given variables]
2000-200=m(50) [solve for the unknown variable]
1800/50=m
36=m
.
y=mx+b
y=36x+200 [plug-in the necessary variables]